Question 1 5 Use the example in the video tutorial for the bond duration of a 10-year semi-annual coupon bond, what is the Macalay Duration of the bond if the coupon rate is 5% and YTM is 9%?

Group of answer choices

7.448

6.84

7.895

7.59

Question 2 Using the same bond maturity, now change the coupon rate to 3% and YTM to 5%. What is Macalay Duration?

Group of answer choices

8.571

8.163

7.762

7.392

Answer both if possible but the first person I asked got question 1 wrong!
Answer #1
Question 1
Semi annual coupon = 1000*(5%/2) = 25
Years Cash flow Present value formula Present value Duration D = (PV*T)
0.5 25 25/(1+9%)^0.5 23.95 11.97
1.0 25 25/(1+9%)^1 22.94 22.94
1.5 25 25/(1+9%)^1.5 21.97 32.95
2.0 25 25/(1+9%)^2 21.04 42.08
2.5 25 25/(1+9%)^2.5 20.15 50.39
3.0 25 25/(1+9%)^3 19.30 57.91
3.5 25 25/(1+9%)^3.5 18.49 64.72
4.0 25 25/(1+9%)^4 17.71 70.84
4.5 25 25/(1+9%)^4.5 16.96 76.34
5.0 25 25/(1+9%)^5 16.25 81.24
5.5 25 25/(1+9%)^5.5 15.56 85.60
6.0 25 25/(1+9%)^6.0 14.91 89.44
6.5 25 25/(1+9%)^6.5 14.28 92.81
7.0 25 25/(1+9%)^7.0 13.68 95.73
7.5 25 25/(1+9%)^7.5 13.10 98.24
8.0 25 25/(1+9%)^8.0 12.55 100.37
8.5 25 25/(1+9%)^8.5 12.02 102.15
9.0 25 25/(1+9%)^9.0 11.51 103.60
9.5 25 25/(1+9%)^9.5 11.03 104.74
10.0 1025 1025/(1+9%)^10.0 432.97 4329.71
Total 750.36 5713.77
Macaulay Duration = Duration D / Present value of cash flows
Macaulay Duration = 5713.77/ 750.36
Macaulay Duration = 7.615
Answer : 7.59
Question 2
Semi annual coupon = 1000*(3%/2) = 15
Years Cash flow Present value formula Present value Duration D = (PV*T)
0.5 15 15/(1+5%)^0.5 14.64 7.32
1.0 15 15/(1+5%)^1 14.29 14.29
1.5 15 15/(1+5%)^1.5 13.94 20.91
2.0 15 15/(1+5%)^2 13.61 27.21
2.5 15 15/(1+5%)^2.5 13.28 33.19
3.0 15 15/(1+5%)^3 12.96 38.87
3.5 15 15/(1+5%)^3.5 12.65 44.26
4.0 15 15/(1+5%)^4 12.34 49.36
4.5 15 15/(1+5%)^4.5 12.04 54.19
5.0 15 15/(1+5%)^5 11.75 58.76
5.5 15 15/(1+5%)^5.5 11.47 63.08
6.0 15 15/(1+5%)^6.0 11.19 67.16
6.5 15 15/(1+5%)^6.5 10.92 71.00
7.0 15 15/(1+5%)^7.0 10.66 74.62
7.5 15 15/(1+5%)^7.5 10.40 78.02
8.0 15 15/(1+5%)^8.0 10.15 81.22
8.5 15 15/(1+5%)^8.5 9.91 84.22
9.0 15 15/(1+5%)^9.0 9.67 87.02
9.5 15 15/(1+5%)^9.5 9.44 89.64
10.0 1015 1015/(1+5%)^10.0 623.12 6231.22
Total 848.43 7275.59
Macaulay Duration = Duration D / Present value of cash flows
Macaulay Duration = 7275.59 / 848.43 8.575
Macaulay Duration = 8.575
Answer : 8.571